aio 6. In an arithmetic sequence, = 28 and ans = 223. what is the sum of the first loo terms? (5pts)
Answer #1

that Given 910 = 28 and 275 = 223 an = a + (n-1) ii a10 = 0,+(10-1)d 28 - 9, +9d u 075= 9,+(15-1)d 223 = 9,7740 2 Subtract frSum of first 100 terms (201) + (100-1)3] S10o = loo 50 [2+99(3)] 50 [27297] 50 [299 Sioo = 14,950